OpenStudy (anonymous):

2y+x=6 write the equation of each line in slope intercerp form and idintify the slope

OpenStudy (anonymous):

2y = -x + 6 y = (-1/2) x + 3 the slope = -1/2

OpenStudy (anonymous):

ok with that natasha?

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

no probs plz close the question by clicking on good answer